# On the Approximation Power of SiLU Networks: Exponential Rates and Depth Efficiency

### v1

@@ 3.3 Fast approximation of continuous functions and general smooth functions by SiLU Neural Networks
$\mathcal{F}_{N}=\bigg\{\sum_{i=1}^{N}c_{i}\,\phi_{\alpha_{i},\beta_{i},\tau,\kappa}(x)\ \bigg|\ c_{i},\alpha_{i},\beta_{i}\in\mathbb{R},\ \tau>0,\ \kappa>0,\beta_{i}>\alpha_{i}\bigg\},\ \text{and}\ \mathcal{F}=\bigcup_{N\geq 1}\mathcal{F}_{N},$
such that
@@
$\phi_{a,b,\tau,\kappa}(x)$ is a SiLU-activated neural network, with $2$ hidden layers, width $2$ with total number of parameters $12$.
@@
For all $x\in\mathbb{R}$, we have this limit
$\lim_{\kappa\to\infty}\frac{\text{SiLU}(\kappa x)}{\kappa}=\max(0,x)$ (8)
and the convergence is uniform on every compact subset of $\mathbb{R}$. We now give a simple proof of this fact.
Recall that
$f_{\kappa}(x)=\frac{\mathrm{SiLU}(\kappa x)}{\kappa}=x\,\sigma(\kappa x),\qquad(x)_{+}=\max(0,x).$
We treat separately the cases $x\geq 0$ and $x\leq 0$ to show that for every $R>0$,
$\sup_{|x|\leq R}\bigl|f_{\kappa}(x)-g(x)\bigr|\xrightarrow[\kappa\to\infty]{}0.$
Case $x\geq 0$.
In this region,
$|f_{\kappa}(x)-(x)_{+}|=|x\,(1-\sigma(\kappa x))|.$
Let $y=\kappa x$, so that $y\geq 0$ and $x=y/\kappa$. Then
$x(1-\sigma(\kappa x))=\frac{1}{\kappa}\,y\bigl(1-\sigma(y)\bigr).$
For $y\geq 0$ we have
$1-\sigma(y)=\frac{1}{1+e^{y}}\leq e^{-y},$
hence
$y\bigl(1-\sigma(y)\bigr)\leq ye^{-y}.$
The function $y\mapsto ye^{-y}$ is bounded on $0,\infty)$; it attains its maximum at $y=1$ with value $1/e$. Thus there is a constant $M\leq 1/e$ such that
$y\bigl(1-\sigma(y)\bigr)\leq M\quad\text{for all }y\geq 0.$
Therefore, for all $x\geq 0$,
$|f_{\kappa}(x)-(x)_{+}|\leq\frac{M}{\kappa}.$
Case $x\leq 0$.
Here, $(x)_{+}=0$ and
$|f_{\kappa}(x)-(x)_{+}|=|x|\,\sigma(\kappa x).$
Let $y=-\kappa x\geq 0$. Then
$|x|\,\sigma(\kappa x)=\frac{y}{\kappa}\,\sigma(-y)=\frac{1}{\kappa}\,\frac{y}{1+e^{y}}\leq\frac{1}{\kappa}\,ye^{-y}\leq\frac{M}{\kappa}.$
For every $R>0$ and all $x$ with $|x|\leq R$, we have
$|f_{\kappa}(x)-(x)_{+}|\leq\frac{M}{\kappa},$
hence
$\sup_{|x|\leq R}|f_{\kappa}(x)-(x)_{+}|\leq\frac{M}{\kappa}\xrightarrow[\kappa\to\infty]{}0.$
Thus the convergence is uniform on every compact set. Using the explicit bound $M=1/e$, we obtain the quantitative estimate
$\sup_{|x|\leq R}\bigl|f_{\kappa}(x)-(x)_{+}\bigr|\leq\frac{1}{e\,\kappa}\leq\eta\implies\kappa=\mathcal{O}\left(\frac{1}{\eta}\right).$
So we have
$\frac{\mathrm{SiLU}\left(\kappa(-\frac{x}{\tau})\right)}{\kappa}\xrightarrow[\kappa\to+\infty,\ \tau\to 0^{+}]{}\begin{cases}0&0\leq x\\ +\infty&0>x\end{cases}.$
$1-\frac{2}{\kappa}\text{SiLU}\left(\frac{\kappa(a-x)}{\tau}\right)\xrightarrow[\kappa\to+\infty,\ \tau\to 0^{+}]{}\begin{cases}1&a\leq x\\ -\infty&a>x\end{cases}.$
Subsequently, relation (8) leads us to
$\phi^{+}_{a,\tau,\kappa}(x)\xrightarrow[\kappa\to+\infty,\ \tau\to 0^{+}]{}\begin{cases}1&a\leq x\\ 0&a>x\end{cases}=\mathbf{1}_{[a,+\infty)}(x).$
Consequently [^1] There exist $\tau_{0}$ and $\kappa_{0}$ two strictly positive real numbers such that for all $\tau_{0}>\tau>0$ and $\kappa>\kappa_{0}$, such that $\phi^{+}_{a,\tau,\kappa}(x)-\phi^{+}_{b,\tau,\kappa}(x)\xrightarrow[]{}\mathbf{1}_{[a,+\infty)}(x)-\mathbf{1}_{[b,+\infty)}(x)=\mathbf{1}_{[a,b]}(x).$ ,
$\phi^{+}_{a,\tau,\kappa}(x)-\phi^{+}_{b,\tau,\kappa}(x)\xrightarrow[\kappa\to+\infty,\ \tau\to 0^{+}]{}\mathbf{1}_{[a,+\infty)}(x)-\mathbf{1}_{[b,+\infty)}(x)=\mathbf{1}_{[a,b]}(x).$
The proposition is proven. ∎
@@
$\left\|\psi-\sum_{i=1}^{m}c_{i}\phi_{\alpha_{i},\beta_{i},\tau,\kappa}\right\|_{\infty}\leq\sum_{i=1}^{m}\max_{j}\{|d_{j}|,|c_{j}|\}\,\|\mathbf{1}_{I_{i}}-\phi_{\alpha_{i},\beta_{i},\tau,\kappa}\|_{\infty}<m\cdot\max_{j}\{|d_{j}|,|c_{j}|\}\cdot\delta\leq\eta.$
@@
Let $f\in C(K)$, $\epsilon>0$. Since $K$ is compact and $f$ is continuous, $f$ is uniformly continuous on $K$.
@@
$\forall x,y\in K,\ |x-y|<\eta\Rightarrow|f(x)-f(y)|<\frac{\epsilon}{2}$
@@
$|f(x)-\psi(x)|=|f(x)-f(x_{j})|<\frac{\epsilon}{2}$
If we set $\delta=\frac{\epsilon}{4m\cdot\max_{j}\{|d_{j}|,|c_{j}|\}}$, from the triangle inequality we deduce the approximation error of the continuous function $f$
$\|f-\sum_{i=1}^{N}c_{i}\,\phi_{\alpha_{i},\beta_{i},\kappa,\tau}\|_{\infty}\leq\|f-\psi\|_{\infty}+\|\psi-\sum_{i=1}^{N}c_{i}\,\phi_{\alpha_{i},\beta_{i},\kappa,\tau}\|_{\infty}\leq\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon$
@@
For $f\in C([a,b])$ with modulus of continuity $\omega_{f}$, and for $\epsilon>0$, there exists a SiLU-activated neural network, with $2$ hidden layers and width $2N$
$f_{\epsilon}(x)=\sum_{i=1}^{N}c_{i}\phi_{\alpha_{i},\beta_{i},\tau,\kappa}(x)$
@@
$N=\mathcal{O}\left(\frac{b-a}{\omega_{f}^{-1}(\epsilon)}\right),\quad\|f-f_{\epsilon}\|_{\infty}<\epsilon$
The total number of parameters is $13N+1$.
@@
$\mathcal{N}_{M,k}$ is a RNN with SiLU activation like in Corollary 2
@@
$\sup_{|x|\leq B}|f(x)-\mathcal{N}_{M,k}(x)|$ $=\sup_{|x|\leq B}|f(x)-P_{f}(x)+P_{f}(x)-\mathcal{N}_{M,k}(x)|$
$\leq\delta+C_{M}\cdot\max_{m}|a_{m}|\cdot\omega_{M}^{-2k}$
@@
According to Corollary 2, $\mathcal{N}_{M,k}$ can be a RNN with SiLU activation.
@@
Let $\Omega=[-B,B]^{d}$ be a bounded domain. For a multi-index $\alpha=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{N}^{d}$, denote $|\alpha|=\alpha_{1}+\cdots+\alpha_{d}$ and
$D^{\alpha}f=\frac{\partial^{|\alpha|}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}}.$
The Sobolev space $W^{n,\infty}(\Omega)$ is defined by
[…]
@@
$\|f\|_{W^{n,\infty}(\Omega)}=\max_{|\alpha|\leq n}\|D^{\alpha}f\|_{L^{\infty}(\Omega)},$
@@
For any $\varepsilon\in(0,1)$, there exists a neural network $\widetilde{f}_{\varepsilon}$ with SiLU activation function such that:

### v2

@@ 3.5 Approximation of Continuous Functions
$\mathcal{F}_{N}=\bigg\{\sum_{i=1}^{N}c_{i}\,\phi_{\alpha_{i},\beta_{i},\tau,\kappa}(x)\ \bigg|\ c_{i}\in\mathbb{R},\beta_{i}-\alpha_{i},\ \tau,\ \kappa\in\mathbb{R}_{+}\bigg\},$
and $\mathcal{F}=\bigcup_{N\geq 1}\mathcal{F}_{N}$, such that
@@
$\phi_{a,b,\tau,\kappa}(x)$ is a SiLU-activated neural network, with $2$ hidden layers, width $2$ with total number of parameters $11$.
@@
Define $f_{\kappa}(x)=\kappa^{-1}\mathrm{SiLU}(\kappa x)=x\sigma(\kappa x)$. We show uniform convergence to $(x)_{+}=\max(0,x)$ on compacts.
For $x\geq 0$: $|f_{\kappa}(x)-x|=x(1-\sigma(\kappa x))\leq\frac{ye^{-y}}{\kappa}\leq\frac{1}{e\kappa}$ with $y=\kappa x$.
For $x\leq 0$: $|f_{\kappa}(x)|=|x|\sigma(\kappa x)\leq\frac{ye^{-y}}{\kappa}\leq\frac{1}{e\kappa}$ with $y=-\kappa x$.
Hence $\sup_{|x|\leq R}|f_{\kappa}(x)-(x)_{+}|\leq(e\kappa)^{-1}\to 0$ as $\kappa\to\infty$.
Setting $f_{\kappa}(-x/\tau)$ gives a pointwise approximation of the Heaviside step:
$\phi_{a,\tau,\kappa}^{+}(x)=1-\frac{2}{\kappa}\mathrm{SiLU}\left(\frac{\kappa(a-x)}{\tau}\right)\xrightarrow[\kappa\to\infty,\tau\to 0^{+}]{}\begin{cases}1,\ a\leq x\\ 0,\ a>x\end{cases}$
i.e., $\phi_{a,\tau,\kappa}^{+}\to\mathbf{1}_{[a,\infty)}$. Therefore
$\phi_{a,\tau,\kappa}^{+}(x)-\phi_{b,\tau,\kappa}^{+}(x)\xrightarrow[\kappa\to\infty,\tau\to 0^{+}]{}\mathbf{1}_{[a,b]}(x),$
which is exactly the claimed limit. ∎
The theoretical convergence shown above is confirmed numerically in Fig. 9, which displays the approximation of $\mathbf{1}_{[1,4]}(x)$ using $\phi^{+}_{1,\tau,\kappa}(x)-\phi^{+}_{4,\tau,\kappa}(x)$.
@@
$\Bigg\|\psi$- $\sum_{i=1}^{m}c_{i}\phi_{\alpha_{i},\beta_{i},\tau,\kappa}\Bigg\|_{\infty}$
$\leq\sum_{i=1}^{m}\max_{j}\{|d_{j}|,|c_{j}|\}\,\|\mathbf{1}_{I_{i}}-\phi_{\alpha_{i},\beta_{i},\tau,\kappa}\|_{\infty}$
$\leq m\cdot\max_{j}\{|d_{j}|,|c_{j}|\}\cdot\delta$
$\leq\eta.$
@@
Let $f\in C(K)$, $\varepsilon>0$. Since $K$ is compact and $f$ is continuous, $f$ is uniformly continuous on $K$.
@@
$\forall x,y\in K,\ |x-y|<\eta\Rightarrow|f(x)-f(y)|<\frac{\varepsilon}{2}$
@@
$|f(x)-\psi(x)|=|f(x)-f(x_{j})|<\frac{\varepsilon}{2}$
If we set $\delta=\frac{\varepsilon}{4m\cdot\max_{j}\{|d_{j}|,|c_{j}|\}}$, we deduce the approximation error of the continuous function $f$
$\|f-\sum_{i=1}^{N}c_{i}\,\phi_{\alpha_{i},\beta_{i},\kappa,\tau}\|_{\infty}\leq\|f-\psi\|_{\infty}+\|\psi-\sum_{i=1}^{N}c_{i}\,\phi_{\alpha_{i},\beta_{i},\kappa,\tau}\|_{\infty}\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon$
@@
For $f\in C([a,b])$ with modulus of continuity $\omega_{f}$, and for $\varepsilon>0$, there exists a SiLU-activated neural network, with $2$ hidden layers and width $2N$
$f_{\varepsilon}(x)=\sum_{i=1}^{N}c_{i}\phi_{\alpha_{i},\beta_{i},\tau,\kappa}(x)$
@@
$N=\mathcal{O}\left(\frac{b-a}{\omega_{f}^{-1}(\varepsilon)}\right),\quad\|f-f_{\varepsilon}\|_{\infty}<\varepsilon$
The total number of parameters is $10N+1$.
The second approach leverages the classical Weierstrass approximation theorem combined with our efficient polynomial implementation (Theorem 4).
@@
$\mathcal{N}_{M,k}$ is a NN with SiLU activation like in Theorem 4.
@@
$|f(x)-\mathcal{N}_{M,k}(x)|\leq|f(x)-P_{f}(x)|+|P_{f}(x)-\mathcal{N}_{M,k}(x)|$
$\sup_{|x|\leq B}|f(x)-\mathcal{N}_{M,k}(x)|\leq\delta+C_{M}\cdot\max_{m}|a_{m}|\cdot\omega_{M}^{-2k}$
@@
Let $\Omega=[-B,B]^{d}$, with $B\leq 1$, be a bounded domain. In this section, vectors are denoted in bold. For a multi-index $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{N}^{d}$, denote $|\bm{\alpha}|=\alpha_{1}+\cdots+\alpha_{d}$ and $D^{\bm{\alpha}}f$ a weak derivative of $f.$ The Sobolev space $W^{n,\infty}(\Omega)$ is defined by
$W^{n,\infty}(\Omega)=\big\{f\in L^{\infty}(\Omega)\;\big|\forall\bm{\alpha}:\ |\bm{\alpha}|\leq n,D^{\bm{\alpha}}f\in L^{\infty}(\Omega)\}.$
@@
$\|f\|_{W^{n,\infty}(\Omega)}=\max_{|\bm{\alpha}|\leq n}\|D^{\bm{\alpha}}f\|_{L^{\infty}(\Omega)},$
@@
For any $f\in\mathcal{F}_{n,d}$ and $\varepsilon\in(0,1)$, there exists a neural network $\widetilde{f}_{\varepsilon}$ with SiLU activation function such that: